• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name eschulte

    Log entry time 00:01:23 on August 02, 2006

    Entry number 176740

    keyword=Swing Shift Log (minor correction)

    16:00-18:45 - Moller, spin dance
    18:58 Getting ready to take data: Wien angle 17.22 Half-wave plate in.
    MCC finishing up from Spin Dance (Mott measurement & restoring vertical
    polarization). This will take about 2 hours. Expect beam back around
    21:00. We will return with 15cm LD2, 4% radiator IN, and approx, 10 uA
    for about 11 M events. Restoration run number: 1985 (HRS-L).

    19:08 HRS-L Q2 POrb alarm, which resolved itself.
    19:22 HRS-L Q2 POrb alarm, which resolved itself.

    20:42 HRS-L Q3 current drifitng causing "red" field alarms. These alarms
    resolve themselves. Field drifts to 255.94 A up from a nominal 255.8
    amps or there about.

    20:46 After a few minutes the HRS-L Q3 current supply appears to have
    "re-regulated" itself, and the current dropped back down to about 255.6
    A, giving a momentum of 0.6419 GeV/c or so. We didn't do anything to
    affect this change.

    20:54 HRS-L Q2 POrb alarm, which resolved itself.

    21:15 Called MCC says that they are getting the machine set up to restore
    beam. The current estimate is that beam will be back in about 1/2 hour.

    22:17 One hour into our half-hour restart!

    22:41 HRS-L Q2 POrb alarm, which resolved itself.


    23:56 At end of shift beam has not yet been restored to Hall A, so the
    Owl shift will have to complete KIN #6 (I hope!).